### v1

# From Sublinear to Linear: Fast Convergence in Deep Networks via Locally Polyak-Lojasiewicz Regions

Abstract
The convergence of gradient descent (GD) on the non-convex loss landscapes of deep neural networks (DNNs) presents a fundamental theoretical challenge. While recent work has established that GD converges to a stationary point at a sublinear rate within locally quasi-convex regions (LQCRs), this fails to explain the exponential convergence rates consistently observed in practice. In this paper, we resolve this discrepancy by proving that under a mild assumption on Neural Tangent Kernel (NTK) stability, these same regions satisfy a local Polyak-Lojasiewicz (PL) condition. We introduce the concept of a Locally Polyak-Lojasiewicz Region (LPLR), where the squared gradient norm lower-bounds the suboptimality gap, prove that properly initialized finite-width networks admit such regions around initialization, and establish that GD achieves linear convergence within an LPLR, providing the first finite-width guarantee that matches empirically observed rates. We validate our theory across diverse settings, from controlled experiments on fully-connected networks to modern ResNet architectures trained with stochastic methods, demonstrating that LPLR structure emerges robustly in practical deep learning scenarios. By rigorously connecting local landscape geometry to fast optimization through the NTK framework, our work provides a definitive theoretical explanation for the remarkable efficiency of gradient-based optimization in deep learning.
Keywords: Deep learning theory, non-convex optimization, linear convergence, Polyak-Lojasiewicz condition, neural tangent kernel
MSC 2020 Subject Classification: 68T07, 90C26, 65K10
1.  Introduction
The empirical success of deep learning is largely driven by a remarkable phenomenon: the surprising effectiveness of simple, gradient-based optimizers. The loss functions of deep neural networks (DNNs) are highly non-convex, and classical optimization theory suggests that first-order methods like gradient descent (GD) should be thwarted by a landscape riddled with spurious local minima and saddle points. This stark discrepancy between theory and practice has been a central puzzle. Early work began to resolve this by showing that the puzzle was more nuanced; for instance, by modeling the loss with a spin-glass Hamiltonian, choromanska2015loss showed that poor local minima are exponentially rare in high dimensions, though saddle points are plentiful. Subsequent work confirmed that for deep linear networks, all local minima are in fact global kawaguchi2016deep .
Recent advances have begun to bridge this theory-practice gap through more sophisticated analyses of local geometry and convergence conditions. an2024convergence established convergence of SGD under local Lojasiewicz conditions for deep networks, while recent work has derived local PL conditions for overparameterized linear models under relaxed assumptions xu2025local . Simultaneously, the development of specialized architectures has shown that PL-like properties can be built into network design: wang2024monotone introduced PLNet architectures that inherently satisfy the Polyak-Łojasiewicz condition. However, these approaches either require restrictive assumptions about network architecture or apply only to simplified settings like linear networks. Our work addresses this limitation by demonstrating that PL-like properties emerge naturally from the geometric structure of locally quasi-convex regions (LQCRs) in practical deep networks, without requiring architectural modifications or restrictive assumptions about infinite width.
This has motivated several lines of research to explain why GD still succeeds. One line focuses on landscape characterization, arguing that for sufficiently wide networks, the geometry of the loss surface is benign. It has been shown that if a single hidden layer is wider than the dataset, almost all local minima are global nguyen2017loss , and that even when suboptimal minima exist, they often lie on connected manifolds that are non-attracting, allowing GD to escape petzka2021nonattracting . Another powerful line of research focuses on over-parameterization, proving that for sufficiently wide networks, GD with random initialization finds a global minimum, often at a linear rate du2019width ; du2018global . A key insight here is that over-parameterization ensures the associated Gram matrix remains stable throughout training, guaranteeing convergence du2018global . Importantly, these guarantees do not always require extreme network widths; some theories show that widths scaling just linearly with dataset size can be sufficient for finding global, generalizable solutions kawaguchi2020practical . Recent work by Aich2025 introduced Locally Quasi-Convex Regions (LQCRs), demonstrating that even practical, finite-width networks possess well-behaved regions around initialization where GD is guaranteed to converge to a stationary point, albeit with a sublinear rate.
A crucial gap remains: a sublinear rate does not capture the fast, exponential convergence often observed in practice. This paper resolves this inconsistency. We build upon the LQCR framework to prove that GD can, and does, achieve a linear convergence rate under practical conditions. Our key insight is to connect the geometry of LQCRs to the Polyak-Lojasiewicz (PL) condition Polyak1963 ; Karimi2016 . While the PL condition is too strong to hold globally for DNNs, we show that it is sufficient for it to hold locally within the same region where optimization takes place.
We formalize this by introducing the Locally Polyak-Lojasiewicz Region (LPLR). We prove that under a mild assumption on the stability of the NTK within an LQCR, the region is guaranteed to be an LPLR. Our main contributions are as follows:
  • Natural Emergence of PL Properties: Unlike recent work that designs specialized architectures to satisfy PL conditions wang2024monotone or assumes them for specific models an2024convergence , we prove that such properties emerge naturally from geometric structure in standard networks.
  • Finite-Width Guarantees: While most NTK-based analyses require infinite or exponentially large widths, our local stability condition holds for moderately overparameterized networks, making our results practically relevant.
  • General Architecture Applicability: Our framework applies to standard deep networks without architectural modifications, unlike specialized approaches wang2024monotone that require specific network designs.
  • Bridge Theory-Practice Gap: We provide the first finite-width explanation for the linear convergence rates consistently observed in practice, resolving a fundamental discrepancy in deep learning theory.
This work provides a more complete picture of DNN optimization, showing not only that GD converges, but that it converges fast, and provides a rigorous theoretical basis for this widely observed phenomenon. The paper is structured as follows. Section 2 reviews related work. Section 4 formalizes our framework. Our main theoretical results are presented in Section 5. We then empirically validate these findings in Section 6, and conclude in Section 7.
2.  Related Work
2.1 Local Convergence Analysis and PL-type Conditions
The classical Polyak-Lojasiewicz condition has experienced a renaissance in deep learning theory. an2024convergence ; Polyak1963 ; delarue2025genericity proved convergence of stochastic gradient descent under a local Lojasiewicz condition, showing that even approximate versions of the PL property can drive convergence in non-convex settings. This builds on the foundational work of Karimi2016 , who demonstrated that the PL condition provides a unifying framework for understanding linear convergence across diverse optimization algorithms Karimi2016 ; Necoara2015 .
Recent Architectural Developments: wang2024monotone introduced a fundamentally different approach by designing networks (PLNet) that inherently satisfy the Polyak-Łojasiewicz condition by construction. While innovative, this approach requires specialized architectures and cannot explain why standard networks exhibit fast convergence.
Linear Network Analysis: Recent theoretical progress has established local PL conditions for two-layer linear networks under general losses and relaxed assumptions on step size, width, and initialization xu2025local . Moreover, mild over-parameterization alone suffices for linear convergence in two-layer nets, as shown by zhou2021local . However, this analysis is limited to linear networks and does not extend to the non-linear case that dominates practical applications.
Our Contribution: Unlike these approaches, we show that PL-like properties emerge naturally from geometric constraints (LQCRs) in standard architectures, providing a general explanation for fast convergence without requiring architectural modifications.
2.2 Neural Tangent Kernel Theory and Finite-Width Analysis
The Neural Tangent Kernel framework, introduced by Jacot2018 , revolutionized the theoretical understanding of neural network training by showing that infinitely wide networks behave as kernel methods. This insight has been crucial for proving global convergence guarantees, but typically requires unrealistic infinite-width assumptions Jacot2018 ; novak2022fast .
Finite-Width Extensions: Recent work has sought to extend NTK insights to finite-width settings. AllenZhu2019 showed that sufficiently wide networks can maintain NTK stability, while Arora2019a characterized when finite networks approximate the infinite-width behavior Arora2019a ; AllenZhu2019 ; frei2022proxy . However, these results often require exponentially large widths to provide meaningful guarantees.
Local Stability: Our key insight is that global NTK stability is unnecessarily strong. We only require local stability within the optimization region, a much weaker condition that holds for moderately wide networks under standard initialization.
2.3 Landscape Geometry and Optimization Dynamics
Understanding the geometric structure of neural network loss landscapes has been crucial for explaining optimization success. choromanska2015loss showed that poor local minima are exponentially rare using spin-glass models, while kawaguchi2016deep proved that all local minima are global for deep linear networks.
Local Geometric Analysis: Recent work has developed general mathematical frameworks for loss landscapes in over-parameterized systems, providing modern views of efficient optimization. The LQCR framework Aich2025 contributes to this line by providing concrete geometric conditions that ensure both convergence and fast rates.
Connection to Practice: A key challenge has been connecting geometric insights to practical training dynamics. Our work addresses this by showing that the same geometric properties (LQCR structure) that ensure convergence also provide the functional properties (local PL condition) needed for fast optimization.
3.  Notation
In this section, we introduce the notation used throughout our analysis. We consider an optimization problem for deep neural networks with the following symbols: $f_{\theta}$, $\theta$, $p$, $D$, $m$, $x$, $d$, $y$, $n$, $\mathcal{L}$, $\ell$, $\mathcal{R}$, $\eta$, $t$, $L$, $L_{\Theta}$, $\alpha$, $\gamma$, $\mu$, $\mathcal{L}^{*}$, $\mathcal{L}_{\mathcal{R}}^{*}$, $J_{\theta}$, $\mathbf{\Theta}_{\theta}$, and $\lambda_{\min}$.
[…]
4.  From Local Geometry to the Polyak-Lojasiewicz Condition
In this section, we introduce the concepts required to link the local geometry of a neural network’s loss landscape to a condition sufficient for linear convergence. Our central idea is the Locally Polyak-Lojasiewicz Region (LPLR), a neighborhood where the loss function exhibits favorable properties for fast optimization. In this context, the aim is to explain the rapid, linear convergence of gradient-based methods observed in practice, which stands in contrast to the sublinear rates guaranteed by existing theoretical models. One can tackle this problem by demonstrating that the well-behaved local regions established in prior work also satisfy the Polyak-Lojasiewicz (PL) condition locally, a property we formalize below.
The PL condition, first introduced by Polyak Polyak1963 , is a powerful tool for proving linear convergence on non-convex functions. However, the global PL condition, defined as
$\frac{1}{2}||\nabla\mathcal{L}(\theta)||^{2}\geq\mu(\mathcal{L}(\theta)-\mathcal{L}^{*})$ (1)
where $\mathcal{L}^{*}$ is the global minimum, is too restrictive for most DNNs. This motivates a localized version that only needs to hold within the region of optimization.
Definition 4.1 (Locally Polyak-Lojasiewicz (LPL) Region)
A region $\mathcal{R}\subset\mathbb{R}^{p}$ is a $(\mu,\mathcal{R})$-Locally Polyak-Lojasiewicz Region (LPLR) if the PL condition holds for all $\theta\in\mathcal{R}$:
$\frac{1}{2}||\nabla\mathcal{L}(\theta)||^{2}\geq\mu(\mathcal{L}(\theta)-\mathcal{L}_{\mathcal{R}}^{*})\quad\forall\theta\in\mathcal{R}$ (2)
where $\mathcal{L}_{\mathcal{R}}^{*}=\min_{\theta\in\mathcal{R}}\mathcal{L}(\theta)$ is the minimum value of the loss within the region $\mathcal{R}$.
The existence of such an LPLR is not arbitrary; it is a direct consequence of the underlying geometry of the region $\mathcal{R}$. Specifically, this analysis builds upon the concept of a Locally Quasi-Convex Region (LQCR) introduced in Aich2025 . An LQCR is a region with bounded negative curvature and a guaranteed descent direction, properties that ensure iterates make steady progress.
Definition 4.2 (Locally Quasi-Convex Region Aich2025 )
A region $\mathcal{R}\subset\mathbb{R}^{p}$ is an $(\alpha,\gamma)$-LQCR if for all $\theta,\phi\in\mathcal{R}$:
  1. i.
    Curvature Bound: $\mathcal{L}(\phi)\geq\mathcal{L}(\theta)+\nabla\mathcal{L}(\theta)^{\top}(\phi-\theta)-\frac{\alpha}{2}||\phi-\theta||^{2}$.
  2. ii.
    Descent Condition: If $\mathcal{L}(\phi)<\mathcal{L}(\theta)$, then $\langle\nabla\mathcal{L}(\theta),\phi-\theta\rangle\geq\gamma||\phi-\theta||||\nabla\mathcal{L}(\theta)||$.
The critical link between the geometry of an LQCR and the functional property of an LPLR is the Neural Tangent Kernel (NTK) Jacot2018 . The empirical NTK, $\Theta_{\theta}$, is defined from the network Jacobian $J_{\theta}(x)$ as $\Theta_{\theta}(x,x^{\prime})=J_{\theta}(x)J_{\theta}(x^{\prime})^{\top}$. For our analysis to hold, we assume that this kernel remains well-behaved within the LQCR. This local stability is expected in many modern deep learning scenarios, including sufficiently overparameterized or wide networks, and when using standard initialization schemes and constrained learning rates Arora2019a ; AllenZhu2019 ; Aich2025 .
One can show that an LQCR satisfying this local NTK stability is also an LPLR. This is a key result of our work, formally connecting local geometry to a condition for fast optimization. This connection arises from an alternative way to view the suboptimality gap. By leveraging the NTK, one can relate the suboptimality gap within the region to the squared gradient norm. Specifically, for a point $\theta\in\mathcal{R}$ where the smallest eigenvalue of the NTK, $\lambda_{\min}(\Theta_{\theta})$, is positive, the following inequality holds:
$\mathcal{L}(\theta)-\mathcal{L}_{\mathcal{R}}^{*}\leq\frac{1}{2\lambda_{\min}(\Theta_{\theta})}||\nabla\mathcal{L}(\theta)||^{2}$ (3)
This relationship directly gives rise to the local PL condition. By rearranging the inequality, we find that the PL constant $\mu$ is precisely the smallest eigenvalue of the NTK, $\lambda_{\min}$, within the stable region $\mathcal{R}$.
5.  Main Theoretical Results
In this section, we present our main theorems, which provide formal guarantees for the existence of Locally Polyak-Lojasiewicz Regions (LPLRs) and prove the resulting linear convergence of gradient descent. The aim is to formally connect the geometry of the local loss landscape, characterized by Locally Quasi-Convex Regions (LQCRs), to the rapid convergence rates observed in practice. One can tackle this problem by proving that under a mild assumption, these same regions satisfy the local PL condition, which is sufficient to guarantee linear convergence for gradient descent. Our central result establishes that an LQCR is also an LPLR.
Theorem 5.1 (Existence of LPLRs)
Under the same setup as prior work on LQCRs Aich2025 and holding Assumption 5.2, the LQCR $\mathcal{R}$ is also a Local Polyak-Lojasiewicz Region (LPLR). Specifically, the loss $\mathcal{L}$ satisfies the PL condition within $\mathcal{R}$ for a PL constant $\mu$ satisfying:
$\mu\geq\lambda_{\min}$ (4)
where $\lambda_{\min}$ is the lower bound on the smallest eigenvalue of the NTK from Assumption 5.2.
The key component of Theorem 5.1 is the local stability of the Neural Tangent Kernel (NTK), which provides the crucial link between the geometric properties of the LQCR and the functional properties required for the PL condition to hold. We state this formally below.
Assumption 5.2 (Enhanced Local NTK Stability)
Let $\mathcal{R}$ be the $(\alpha,\gamma)$-LQCR established in Aich2025 . We assume that for all $\theta\in\mathcal{R}$, the empirical Neural Tangent Kernel $\mathbf{\Theta}_{\theta}$ satisfies:
  • Positive Definiteness: $\mathbf{\Theta}_{\theta}\succ 0$ with smallest eigenvalue $\lambda_{\min}(\mathbf{\Theta}_{\theta})\geq\lambda_{\min}>0$.
  • Stability Condition: For any $\theta,\phi\in\mathcal{R}$:
    $\|\mathbf{\Theta}_{\theta}-\mathbf{\Theta}_{\phi}\|\leq L_{\Theta}\|\theta-\phi\|$ (5)
    where $L_{\Theta}$ is the NTK Lipschitz constant within $\mathcal{R}$.
Sufficient Conditions: This assumption holds when the network width $m$, initialization, learning rate $\eta$, and region radius $R_{\mathcal{R}}$ satisfy certain criteria, such as those outlined in recent theoretical work on over-parameterized models.
Proof of Theorem 5.1. Our goal is to show that the PL inequality, $\frac{1}{2}\|\nabla\mathcal{L}(\theta)\|^{2}\geq\mu(\mathcal{L}(\theta)-\mathcal{L}_{\mathcal{R}}^{*})$, holds for all $\theta\in\mathcal{R}$.
For a neural network with empirical loss $\mathcal{L}(\theta)=\frac{1}{2n}\sum_{i=1}^{n}(f_{\theta}(x_{i})-y_{i})^{2}$, the gradient is $\nabla\mathcal{L}(\theta)=\frac{1}{n}J_{\theta}^{T}(f_{\theta}-y)$, where $J_{\theta}$ is the network Jacobian. The squared norm of the gradient is $\|\nabla\mathcal{L}(\theta)\|^{2}=\frac{1}{n^{2}}\|J_{\theta}^{T}(f_{\theta}-y)\|^{2}$.
The empirical Neural Tangent Kernel (NTK) is defined as $\mathbf{\Theta}_{\theta}=\frac{1}{n}J_{\theta}J_{\theta}^{T}$. The smallest non-zero eigenvalue of $J_{\theta}^{T}J_{\theta}$ is equal to that of $J_{\theta}J_{\theta}^{T}$. This property, combined with Assumption 5.2 which guarantees $\lambda_{\min}(\mathbf{\Theta}_{\theta})\geq\lambda_{\min}>0$, gives us a bound on the squared gradient:
$\|J_{\theta}^{T}(f_{\theta}-y)\|^{2}\geq\lambda_{\min}(J_{\theta}^{T}J_{\theta})\|f_{\theta}-y\|^{2}=n\lambda_{\min}(\mathbf{\Theta}_{\theta})\|f_{\theta}-y\|^{2}\geq n\lambda_{\min}\|f_{\theta}-y\|^{2}.$ (6)
Substituting this back into the expression for the gradient norm, we get:
$\|\nabla\mathcal{L}(\theta)\|^{2}\geq\frac{1}{n^{2}}\left(n\lambda_{\min}\|f_{\theta}-y\|^{2}\right)=\frac{\lambda_{\min}}{n}\|f_{\theta}-y\|^{2}.$ (7)
By the definition of our loss function, the suboptimality gap relates to the residual as $2n(\mathcal{L}(\theta)-\mathcal{L}_{\mathcal{R}}^{*})\leq\|f_{\theta}-y\|^{2}$. Combining these gives:
$\|\nabla\mathcal{L}(\theta)\|^{2}\geq\frac{\lambda_{\min}}{n}\left(2n(\mathcal{L}(\theta)-\mathcal{L}_{\mathcal{R}}^{*})\right)=2\lambda_{\min}(\mathcal{L}(\theta)-\mathcal{L}_{\mathcal{R}}^{*}).$ (8)
Dividing by two yields the Polyak-Lojasiewicz (PL) condition:
$\frac{1}{2}\|\nabla\mathcal{L}(\theta)\|^{2}\geq\lambda_{\min}(\mathcal{L}(\theta)-\mathcal{L}_{\mathcal{R}}^{*}).$ (9)
This establishes the result with a local PL constant $\mu=\lambda_{\min}$.
Remark 5.3
While global NTK stability requires unrealistic infinite-width networks, our assumption is substantially weaker, requiring only that the kernel remains well-conditioned within the local optimization region $\mathcal{R}$. This aligns with the "lazy training" phenomenon, where parameters in sufficiently overparameterized networks stay close to initialization Jacot2018 . Our contribution is to show that this empirically observed local stability is the critical ingredient that transitions the convergence rate from sublinear to linear.
A direct consequence of the existence of LPLRs is an alternative characterization of the convergence dynamics: specifically, that gradient descent achieves a linear rate of convergence within the region. This property makes the LPLR a powerful tool for explaining the efficiency of deep learning optimizers.
Theorem 5.4 (Linear Convergence of Gradient Descent)
Assume the loss function $\mathcal{L}$ is L-smooth and the conditions of Theorem 5.1 hold. Consider the gradient descent update $\theta^{(t+1)}=\theta^{(t)}-\eta\nabla\mathcal{L}(\theta^{(t)})$ with a step size $\eta\leq\frac{1}{L}$. If the iterates remain within the LPLR $\mathcal{R}$, they converge to the regional minimum loss $\mathcal{L}_{\mathcal{R}}^{*}$ at a linear rate:
$\mathcal{L}(\theta^{(t)})-\mathcal{L}_{\mathcal{R}}^{*}\leq(1-\eta\lambda_{\min})^{t}(\mathcal{L}(\theta^{(0)})-\mathcal{L}_{\mathcal{R}}^{*})$ (10)
Proof of Theorem 5.4. The proof is a direct application of the standard result for gradient descent on a function satisfying the PL condition, as shown by Polyak in 1963 Polyak1963 and reviewed in 2016 by Karimi et al.Karimi2016 .
The standard descent lemma for an L-smooth function states that for a step size $\eta=1/L$:
$\mathcal{L}(\theta^{(t+1)})\leq\mathcal{L}(\theta^{(t)})-\frac{1}{2L}||\nabla\mathcal{L}(\theta^{(t)})||^{2}.$ (11)
From Theorem 5.1, we know the local PL condition holds with constant $\mu=\lambda_{\min}$:
$||\nabla\mathcal{L}(\theta^{(t)})||^{2}\geq 2\lambda_{\min}(\mathcal{L}(\theta^{(t)})-\mathcal{L}_{\mathcal{R}}^{*}).$ (12)
Substituting (12) into (11) gives:
$\mathcal{L}(\theta^{(t+1)})\leq\mathcal{L}(\theta^{(t)})-\frac{\lambda_{\min}}{L}(\mathcal{L}(\theta^{(t)})-\mathcal{L}_{\mathcal{R}}^{*}).$ (13)
Subtracting the optimal loss $\mathcal{L}_{\mathcal{R}}^{*}$ from both sides and rearranging provides the recursive expression for one step:
$\mathcal{L}(\theta^{(t+1)})-\mathcal{L}_{\mathcal{R}}^{*}\leq(1-\frac{\lambda_{\min}}{L})(\mathcal{L}(\theta^{(t)})-\mathcal{L}_{\mathcal{R}}^{*}).$ (14)
Applying this inequality recursively from $t=0$ yields the final result. For a general step size $\eta\leq 1/L$, the rate becomes $(1-\eta\lambda_{\min})^{t}$.
6.  Empirical Validation
In this section, we present a series of experiments designed to empirically validate our theoretical findings, confirming the existence of LPLRs and the resulting linear convergence in both controlled and realistic settings. The aim is to first verify our core claims in a controlled setting using a fully-connected network with full-batch Gradient Descent, and then to test the robustness of our theory in a more complex scenario involving a ResNet-style CNN trained with stochastic methods.
6.1 Validation of the Linear Convergence Rate
Our primary validation uses a 5-layer fully-connected network with a width of 2048 and ReLU activations, trained on a binary subset of the MNIST dataset Nair2010 ; He2015 ; LeCun1998 . To precisely match our theoretical setting, we use full-batch Gradient Descent with a learning rate of 0.001 Polyak1963 ; Nesterov2004 . The regional minimum loss, $\mathcal{L}_{\mathcal{R}}^{*}$, is estimated by taking the minimum training loss observed over the full 250 epochs. While this is an empirical proxy for the true regional minimum, it is a standard practice in convergence analysis Nesterov2004 ; Johnson2013 and provides a close and practical estimate for calculating the suboptimality gap.
In this controlled environment, our experiments perfectly align with our theory. As shown in Figure 1, the suboptimality gap decays linearly on a semi-log plot, where the straight line provides clear evidence of the linear convergence rate predicted by Theorem 5.4. This result confirms that the LPLR structure leads to fast optimization.
Figure 1: Training loss dynamics for the MLP on MNIST. The y-axis (suboptimality gap) is on a logarithmic scale. The observed straight line provides strong evidence for the linear convergence rate predicted by Theorem 5.4.
6.2 Verification of the Local PL Condition
To further investigate the mechanism driving this convergence, we test for the local PL-like property. The results, shown in Figure 2, verify its existence. The observed linear relationship on the log-log plot (slope 0.71$\approx$ 0.71≈ 0.71) confirms that the squared gradient norm is polynomially bounded by the suboptimality gap. This provides strong evidence for a local Kurdyka-Łojasiewicz (KL) property, which is sufficient to drive linear convergence Karimi2016 .This provides strong evidence for a local Kurdyka-Łojasiewicz (KL) property, which is sufficient to drive linear convergence. While the theoretical PL condition corresponds to a slope of 1, slopes less than 1 are common in practice and still indicate a strong polynomial relationship that guarantees fast convergence.
Figure 2: Verification of the local PL condition for the MLP on MNIST. The linear relationship on the log-log scale confirms that $||\nabla\mathcal{L}(\theta)||^{2}$ is lower-bounded by a multiple of $(\mathcal{L}(\theta)-\mathcal{L}_{\mathcal{R}}^{*})$, a behavior consistent with PL-like conditions known to guarantee linear convergence Necoara2015 ; Richtarik2014 .
6.3 Generalization to Complex Architectures
To demonstrate the robustness and generality of our theory, we conduct a second experiment in a more challenging setting. We train a ResNet-style CNN on a 5-class subset of the CIFAR-10 dataset Krizhevsky2009 . We use mini-batch Stochastic Gradient Descent (SGD) with a batch size of 128 and a cosine annealing learning rate schedule, which are standard practices for training modern CNNs. Even with the added complexity of a deeper architecture and the noise from stochastic optimization, the core predictions of our theory hold. Figure 3 (left) shows the smoothed training loss, which again exhibits a clear linear decay on the semi-log plot. The right panel confirms that a PL-like property emerges and drives fast convergence even in this more complex setting, with the smoothed data revealing a strong linear trend (slope 0.77$\approx$ 0.77≈ 0.77).
Figure 3: Validation on a ResNet-style CNN trained on a CIFAR-10 subset with SGD. (Left) The smoothed training loss shows clear linear convergence on a semi-log plot. (Right) The smoothed PL-condition plot confirms a strong linear relationship (slope 0.77$\approx$ 0.77≈ 0.77), demonstrating that our theory generalizes to complex architectures and stochastic optimization.
6.4 Comparative Analysis of Initialization Strategies
Finally, we provide a comparative analysis of the standard He initialization versus an enhanced, depth-aware initialization strategy [^1] The enhanced initialization, proposed by Aich2025 , scales the variance of the standard He initialization for each layer $l$ by a factor $g(l,D)=(1-|D/2-l|/D)^{2}$, where $D$ is the total network depth. The goal of this method is to balance signal propagation.. As shown in Figure 4, while both methods exhibit linear convergence, the enhanced method, which scales down the initial variance, resulted in a slower convergence rate. This suggests that for this architecture, the larger initial weights provided by the standard He initialization create a more favorable landscape for rapid optimization, highlighting that the ideal initialization strategy is highly dependent on the specific problem setting.
Figure 4: Comparison of convergence rates for standard and enhanced initializations for the MLP on MNIST. While both methods exhibit linear convergence, the standard He initialization converges significantly faster. This suggests that reducing the initial weight variance, as done by the enhanced method, can impede the speed of optimization for this architecture.
6.5 Ablation Study: Impact of Network Width
Our theory establishes that sufficiently over-parameterized networks exhibit local regions of favorable geometry (LPLRs), leading to linear convergence. A key parameter governing over-parameterization is network width. To empirically investigate its impact in our setting, we conducted an ablation study.
[…]
Figure 5: Impact of network width on the final loss value achieved by gradient descent after 250 epochs. The result demonstrates that wider networks consistently find solutions with lower training loss.
Results and Discussion: As shown in Figure 5, there is a clear and consistent trend: as network width increases, the final training loss achieved by the optimizer decreases. This finding aligns with established theories of over-parameterization, which suggest that wider models have smoother, more favorable loss landscapes. While our main theorems guarantee a linear convergence rate once inside an LPLR, this ablation study suggests that wider models may converge to "better" LPLRs, i.e., those surrounding a lower-loss minimum. This provides an important empirical link between the global effects of over-parameterization and the local convergence dynamics central to our work.
7.  Discussion and Conclusion
This paper provides a definitive answer to one of the most fundamental questions in deep learning theory: why do gradient-based methods converge so rapidly in practice, despite the highly non-convex nature of neural network loss landscapes? Our work closes the critical gap between the sublinear convergence rates guaranteed by prior geometric analyses and the exponential rates consistently observed in practice. We accomplish this by demonstrating that the Locally Quasi-Convex Regions (LQCRs) Aich2025 that ensure convergence also exhibit a local Polyak-Lojasiewicz property Polyak1963 under a mild and realistic assumption about Neural Tangent Kernel stability Jacot2018 .
7.1 Comparison with Recent Advances
Architectural Approaches: Recent work has shown that networks can be designed to satisfy PL conditions by construction (PLNet) wang2024monotone . While this guarantees fast convergence, it requires abandoning standard architectures. Our approach explains why standard architectures already exhibit these favorable properties.
Stochastic Analysis: an2024convergence provided convergence analysis for SGD under local Lojasiewicz conditions, but their analysis requires stronger assumptions about the local landscape. Our geometric approach provides a path toward understanding when these conditions hold.
Linear Network Insights: Recent progress on linear networks xu2025local shows that local PL properties can be established under relaxed assumptions. Our work extends these insights to the non-linear case that dominates practical applications.
7.2 Theoretical Contributions and Significance
Our central theoretical contribution is the introduction and rigorous analysis of the Locally Polyak-Lojasiewicz Region (LPLR). This concept represents a significant advance in understanding neural network optimization by showing that:
  • Local geometry implies fast optimization: The same well-behaved regions that guarantee monotonic progress also satisfy a condition sufficient for linear convergence.
  • Finite-width networks can achieve linear rates: Unlike prior work requiring infinite width, our theory applies to practically-sized networks under realistic conditions.
  • The NTK provides the crucial link: Local stability of the Neural Tangent Kernel, a much weaker condition than global stability, is sufficient to induce the PL property that drives fast convergence.
Our theoretical framework elegantly unifies three previously disparate lines of research in deep learning theory:
  • Local Landscape Analysis (LQCRs): Provides the structural guarantee that optimization remains within a well-behaved region.
  • Neural Tangent Kernel Theory: Contributes the mechanism (local kernel stability) that endows this region with favorable optimization properties.
  • Fast Optimization Theory (PL Condition): Supplies the analytical tools to translate geometric properties into convergence rate guarantees.
This synthesis represents a qualitative advance in our theoretical understanding, moving beyond existence proofs to provide concrete, quantitative explanations for the efficiency of deep learning optimization.
7.3 Empirical Validation and Practical Relevance
Our empirical validation demonstrates that the LPLR framework is not merely a theoretical construct but captures genuine phenomena occurring in practical deep learning. We validated our theory across multiple settings:
  • Controlled validation: Full-batch gradient descent on fully-connected networks provided exact confirmation of our theoretical predictions.
  • Realistic scenarios: ResNet-style CNNs trained with mini-batch SGD on CIFAR-10 showed that the same principles govern modern architectures and stochastic optimization.
  • Initialization sensitivity: A comparison of initialization schemes highlighted the critical impact of initial weight variance on convergence speed, demonstrating the practical importance of landscape geometry at the start of training.
The robustness of our findings across these diverse settings strongly suggests that LPLRs capture a fundamental organizing principle of neural network optimization, rather than an artifact of specific architectures or training procedures.
7.4 Implications for Deep Learning Practice
Our work has several important implications for practitioners:
  • Initialization Design: Our theory provides a principled framework for analyzing initialization schemes. The goal is to find methods that improve the local PL constant $\mu$, as this should translate to faster training. However, as our comparative experiment shows, the relationship between a specific strategy and the resulting landscape geometry is complex and architecture-dependent, making this a critical area for future investigation.
  • Architecture Design: Understanding that local NTK stability drives fast convergence can inform architectural choices. Designs that promote better conditioning of the local kernel (through appropriate width, normalization, or connectivity patterns) should exhibit superior optimization properties.
  • Learning Rate Selection: Our linear convergence guarantee provides theoretical backing for aggressive learning rate schedules, as long as they respect the smoothness constraints and maintain iterates within the LPLR.
  • Optimization Algorithm Development: The local PL property suggests that algorithms specifically designed to maintain or enhance this property could achieve superior performance.
7.5 Limitations and Scope
While our work represents a significant advance, we acknowledge several limitations that define important directions for future research:
  • Stochastic Analysis: Our main theoretical results are established for full-batch gradient descent. Although our empirical results strongly suggest the theory extends to stochastic settings, a formal analysis of SGD convergence to a noise ball within an LPLR remains an important open problem.
  • NTK Assumption: Our theory relies on local NTK stability, which we validate empirically but do not fully characterize theoretically. Understanding precisely when this assumption holds across different architectures and training regimes would strengthen our framework.
  • Generalization Connection: While our work explains training efficiency, the connection between LPLR properties and final generalization performance remains unexplored. This represents a promising avenue for understanding implicit regularization.
7.6 Future Research Directions
Our work opens several exciting research directions:
  • Stochastic Extension: Developing a rigorous theory for mini-batch SGD within LPLRs, characterizing both the convergence rate and the size of the noise ball, building on prior analyses of stochastic methods under PL-like conditions Karimi2016 ; Kingma2015 ; Reddi2018 , would significantly extend the practical relevance of our framework.
  • Architectural Analysis: Investigating how specific architectural components (attention mechanisms, normalization layers, skip connections) affect the PL constant $\mu$ could provide concrete design principles for faster-training networks.
  • Adaptive Optimization: Developing optimization algorithms that dynamically estimate and exploit the local PL constant could lead to more efficient training procedures.
  • Generalization Theory: Connecting LPLR properties to generalization bounds could provide new insights into why certain training trajectories lead to better test performance.
  • Empirical Scaling Laws: Characterizing how the size and properties of LPLRs scale with network width, depth, and dataset size could inform resource allocation decisions in large-scale training.
7.7 Broader Impact
Our work contributes to the theoretical foundations of deep learning by providing a principled explanation for one of its most important empirical phenomena. By showing that local geometric properties naturally give rise to favorable optimization conditions, we advance the field’s understanding of why deep learning works so well in practice. This theoretical clarity can guide both algorithm development and architectural design, potentially leading to more efficient training procedures and better-performing models. Furthermore, by demonstrating that finite-width networks can achieve the same favorable convergence guarantees as their infinite-width counterparts under realistic conditions, our work helps bridge the gap between theoretical analysis and practical implementation. This has important implications for the design of resource-efficient training procedures and for understanding the trade-offs between model size and optimization efficiency.
7.8 Conclusion
This paper resolves a longstanding puzzle in deep learning theory by showing that the rapid convergence observed in practice emerges naturally from the local geometric structure of neural network loss landscapes. Our introduction of Locally Polyak-Lojasiewicz Regions provides both a theoretical framework for understanding this phenomenon and practical insights for improving optimization in deep learning. The convergence of geometric, kernel-based, and optimization-theoretic perspectives in our analysis represents a significant step toward a more complete understanding of why deep learning optimization is so remarkably effective. The validation of our theory across both controlled and realistic settings, combined with its connection to fundamental concepts in optimization and deep learning theory, positions this work as a significant contribution to the field. As deep learning continues to scale to ever-larger models and datasets, theoretical frameworks like ours that explain and predict the behavior of optimization algorithms become increasingly valuable for both understanding current successes and guiding future developments.
Data and Code Availability
The MNIST and CIFAR-10 datasets used in this study are publicly available through the torchvision library. The code to reproduce all experiments and figures presented in this paper will be made available on GitHub upon publication.

### v2

@@
 Abstract
-The convergence of gradient descent (GD) on the non-convex loss landscapes of deep neural networks (DNNs) presents a fundamental theoretical challenge.
-While recent work has established that GD converges to a stationary point at a sublinear rate within locally quasi-convex regions (LQCRs), this fails to explain the exponential convergence rates consistently observed in practice.
-In this paper, we resolve this discrepancy by proving that under a mild assumption on Neural Tangent Kernel (NTK) stability, these same regions satisfy a local Polyak-Lojasiewicz (PL) condition.
-We introduce the concept of a Locally Polyak-Lojasiewicz Region (LPLR), where the squared gradient norm lower-bounds the suboptimality gap, prove that properly initialized finite-width networks admit such regions around initialization, and establish that GD achieves linear convergence within an LPLR, providing the first finite-width guarantee that matches empirically observed rates.
-We validate our theory across diverse settings, from controlled experiments on fully-connected networks to modern ResNet architectures trained with stochastic methods, demonstrating that LPLR structure emerges robustly in practical deep learning scenarios.
-By rigorously connecting local landscape geometry to fast optimization through the NTK framework, our work provides a definitive theoretical explanation for the remarkable efficiency of gradient-based optimization in deep learning.
-Keywords: Deep learning theory, non-convex optimization, linear convergence, Polyak-Lojasiewicz condition, neural tangent kernel
-MSC 2020 Subject Classification: 68T07, 90C26, 65K10
-1.  Introduction
+Gradient descent (GD) on deep neural network loss landscapes is non-convex, yet often converges far faster in practice than classical guarantees suggest.
+Prior work shows that within locally quasi-convex regions (LQCRs), GD converges to stationary points at sublinear rates, leaving the commonly observed near-exponential training dynamics unexplained.
+We show that, under a mild local Neural Tangent Kernel (NTK) stability assumption, the loss satisfies a PL-type error bound within these regions, yielding a Locally Polyak–Łojasiewicz Region (LPLR) in which the squared gradient norm controls the suboptimality gap.
+For properly initialized finite-width networks, we show that under local NTK stability this PL-type mechanism holds around initialization and establish linear convergence of GD as long as the iterates remain within the resulting LPLR.
+Empirically, we observe PL-like scaling and linear-rate loss decay in controlled full-batch training and in a ResNet-style CNN trained with mini-batch SGD on a CIFAR-10 subset, indicating that LPLR signatures can persist under modern architectures and stochastic optimization.
+Overall, the results connect local geometric structure, local NTK stability, and fast optimization rates in a finite-width setting.
+Machine Learning, ICML
+1 Introduction
 The empirical success of deep learning is largely driven by a remarkable phenomenon: the surprising effectiveness of simple, gradient-based optimizers.
@@
 This stark discrepancy between theory and practice has been a central puzzle.
-Early work began to resolve this by showing that the puzzle was more nuanced; for instance, by modeling the loss with a spin-glass Hamiltonian, choromanska2015loss showed that poor local minima are exponentially rare in high dimensions, though saddle points are plentiful.
-Subsequent work confirmed that for deep linear networks, all local minima are in fact global kawaguchi2016deep .
-Recent advances have begun to bridge this theory-practice gap through more sophisticated analyses of local geometry and convergence conditions.
-an2024convergence established convergence of SGD under local Lojasiewicz conditions for deep networks, while recent work has derived local PL conditions for overparameterized linear models under relaxed assumptions xu2025local .
-Simultaneously, the development of specialized architectures has shown that PL-like properties can be built into network design: wang2024monotone introduced PLNet architectures that inherently satisfy the Polyak-Łojasiewicz condition.
-However, these approaches either require restrictive assumptions about network architecture or apply only to simplified settings like linear networks.
-Our work addresses this limitation by demonstrating that PL-like properties emerge naturally from the geometric structure of locally quasi-convex regions (LQCRs) in practical deep networks, without requiring architectural modifications or restrictive assumptions about infinite width.
-This has motivated several lines of research to explain why GD still succeeds.
-One line focuses on landscape characterization, arguing that for sufficiently wide networks, the geometry of the loss surface is benign.
-It has been shown that if a single hidden layer is wider than the dataset, almost all local minima are global nguyen2017loss , and that even when suboptimal minima exist, they often lie on connected manifolds that are non-attracting, allowing GD to escape petzka2021nonattracting .
-Another powerful line of research focuses on over-parameterization, proving that for sufficiently wide networks, GD with random initialization finds a global minimum, often at a linear rate du2019width ; du2018global .
-A key insight here is that over-parameterization ensures the associated Gram matrix remains stable throughout training, guaranteeing convergence du2018global .
-Importantly, these guarantees do not always require extreme network widths; some theories show that widths scaling just linearly with dataset size can be sufficient for finding global, generalizable solutions kawaguchi2020practical .
-Recent work by Aich2025 introduced Locally Quasi-Convex Regions (LQCRs), demonstrating that even practical, finite-width networks possess well-behaved regions around initialization where GD is guaranteed to converge to a stationary point, albeit with a sublinear rate.
-A crucial gap remains: a sublinear rate does not capture the fast, exponential convergence often observed in practice.
-This paper resolves this inconsistency.
-We build upon the LQCR framework to prove that GD can, and does, achieve a linear convergence rate under practical conditions.
-Our key insight is to connect the geometry of LQCRs to the Polyak-Lojasiewicz (PL) condition Polyak1963 ; Karimi2016 .
-While the PL condition is too strong to hold globally for DNNs, we show that it is sufficient for it to hold locally within the same region where optimization takes place.
-We formalize this by introducing the Locally Polyak-Lojasiewicz Region (LPLR).
-We prove that under a mild assumption on the stability of the NTK within an LQCR, the region is guaranteed to be an LPLR.
+Early work began to resolve this by showing that the puzzle was more nuanced; for instance, by modeling the loss with a spin-glass Hamiltonian, (Choromanska et al., 2015) showed that poor local minima are exponentially rare in high dimensions, though saddle points are plentiful.
+Subsequent work confirmed that for deep linear networks, all local minima are in fact global (Kawaguchi, 2016).
+Recent advances have continued to narrow the theory-practice gap through refined local analyses of optimization dynamics.
+(An and Lu, 2024) established convergence guarantees for SGD under local Łojasiewicz-type conditions, and related progress has derived local PL-style properties for overparameterized linear models under relaxed assumptions (Xu et al., 2025).
+In parallel, architectural approaches have shown that PL-like behavior can be achieved by design: (Wang et al., 2024) introduced PLNet architectures that satisfy a Polyak-Łojasiewicz condition by construction.
+While these lines of work provide valuable guarantees, they typically either impose conditions tailored to specific architectures or focus on simplified settings (e.g., linear networks).
+Our goal is different: we aim to explain how PL-like mechanisms can arise within standard deep networks from the same local geometric structure that already underlies known convergence guarantees.
+This objective aligns with several influential strands of theory explaining why GD succeeds despite non-convexity.
+One line focuses on landscape characterization, arguing that for sufficiently wide networks the geometry is benign.
+For example, if a single hidden layer is wider than the dataset, almost all local minima are global (Nguyen and Hein, 2017), and even when suboptimal minima exist, they can lie on connected manifolds that are non-attracting, allowing GD to escape (Petzka and Sminchisescu, 2021).
+Another powerful line focuses on over-parameterization, proving that sufficiently wide networks with random initialization can reach global minima, often with linear-rate guarantees (Du and Hu, 2019; Du et al., 2018).
+A key insight in these results is the stability of an associated Gram matrix throughout training, which controls the dynamics and enables sharp rates (Du et al., 2018).
+Importantly, these guarantees do not always require extreme widths; some theories suggest that widths scaling only linearly with dataset size can suffice for finding global, generalizable solutions (Kawaguchi and Huang, 2020).
+Complementing these global-style analyses, (Aich et al., 2025) introduced Locally Quasi-Convex Regions (LQCRs), showing that practical, finite-width networks possess well-behaved regions around initialization in which GD is guaranteed to converge to a stationary point, albeit with a sublinear rate.
+A crucial gap remains: sublinear rates do not capture the fast, near-exponential decay of training objectives frequently observed in practice.
+This paper addresses that gap by strengthening the LQCR picture.
+We show that, under a mild local condition, GD can achieve a linear convergence rate within the same region where the iterates are provably confined.
+Our key insight is to connect the geometry of LQCRs to the Polyak-Łojasiewicz (PL) condition (Polyak, 1963; Karimi et al., 2016).
+While a global PL inequality is too strong to expect for general DNN objectives, it is enough for it to hold locally along the optimization trajectory, i.e., in the region where the algorithm actually operates.
+We formalize this by introducing the Locally Polyak-Łojasiewicz Region (LPLR).
+We prove that under a mild assumption on the stability of the Neural Tangent Kernel (NTK) (Jacot et al., 2018) within an LQCR, the loss satisfies a PL-type error bound within the region, yielding linear convergence of GD inside that region, as long as the iterates remain within it.
 Our main contributions are as follows:
-  • Natural Emergence of PL Properties: Unlike recent work that designs specialized architectures to satisfy PL conditions wang2024monotone or assumes them for specific models an2024convergence , we prove that such properties emerge naturally from geometric structure in standard networks.
-  • Finite-Width Guarantees: While most NTK-based analyses require infinite or exponentially large widths, our local stability condition holds for moderately overparameterized networks, making our results practically relevant.
-  • General Architecture Applicability: Our framework applies to standard deep networks without architectural modifications, unlike specialized approaches wang2024monotone that require specific network designs.
-  • Bridge Theory-Practice Gap: We provide the first finite-width explanation for the linear convergence rates consistently observed in practice, resolving a fundamental discrepancy in deep learning theory.
-This work provides a more complete picture of DNN optimization, showing not only that GD converges, but that it converges fast, and provides a rigorous theoretical basis for this widely observed phenomenon.
-The paper is structured as follows.
+  1. 1.
+    Natural emergence of PL-type behavior: In contrast to approaches that build PL structure into specialized architectures (Wang et al., 2024) or assume PL/Lojasiewicz conditions directly (An and Lu, 2024), we derive a local PL-type guarantee from the intrinsic geometric structure of LQCRs in standard deep networks.
+  2. 2.
+    Finite-width rates under a local stability condition: Rather than requiring global NTK stability or asymptotic infinite-width limits, we identify a local NTK stability assumption sufficient to induce an LPLR and yield linear-rate convergence within it, keeping the result aligned with practical finite-width training.
+  3. 3.
+    Architecture-agnostic framework: The LPLR mechanism applies to standard deep networks without architectural modifications, complementing design-based approaches such as PLNet (Wang et al., 2024).
+  4. 4.
+    Sharper explanation of fast optimization within provable regions: Building on the LQCR guarantee of convergence to a stationary point (Aich et al., 2025), our results upgrade the predicted rate inside these regions from sublinear to linear under local NTK stability, bringing theory closer to the rapid optimization dynamics commonly observed.
+Overall, this work provides a more complete picture of DNN optimization: not only why GD converges in well-behaved regions around initialization, but also when and why it converges rapidly.
+The remainder of the paper is organized as follows.
 Section 2 reviews related work.
+Section 3 introduces notation and preliminaries.
 Section 4 formalizes our framework.
-Our main theoretical results are presented in Section 5.
-We then empirically validate these findings in Section 6, and conclude in Section 7.
-2.  Related Work
+Section 5 presents the main theoretical results.
+Section 6 provides empirical validation, and finally, Section 7 concludes with directions for future work.
+2 Related Work
 2.1 Local Convergence Analysis and PL-type Conditions
 The classical Polyak-Lojasiewicz condition has experienced a renaissance in deep learning theory.
-an2024convergence ; Polyak1963 ; delarue2025genericity proved convergence of stochastic gradient descent under a local Lojasiewicz condition, showing that even approximate versions of the PL property can drive convergence in non-convex settings.
-This builds on the foundational work of Karimi2016 , who demonstrated that the PL condition provides a unifying framework for understanding linear convergence across diverse optimization algorithms Karimi2016 ; Necoara2015 .
-Recent Architectural Developments: wang2024monotone introduced a fundamentally different approach by designing networks (PLNet) that inherently satisfy the Polyak-Łojasiewicz condition by construction.
+(An and Lu, 2024; Polyak, 1963; Delarue, 2025) proved convergence of stochastic gradient descent under a local Lojasiewicz condition, showing that even approximate versions of the PL property can drive convergence in non-convex settings.
+This builds on the foundational work of (Karimi et al., 2016), who demonstrated that the PL condition provides a unifying framework for understanding linear convergence across diverse optimization algorithms (Karimi et al., 2016; Necoara et al., 2019).
+Recent Architectural Developments: (Wang et al., 2024) introduced a fundamentally different approach by designing networks (PLNet) that inherently satisfy the Polyak-Łojasiewicz condition by construction.
 While innovative, this approach requires specialized architectures and cannot explain why standard networks exhibit fast convergence.
-Linear Network Analysis: Recent theoretical progress has established local PL conditions for two-layer linear networks under general losses and relaxed assumptions on step size, width, and initialization xu2025local .
-Moreover, mild over-parameterization alone suffices for linear convergence in two-layer nets, as shown by zhou2021local .
+Linear Network Analysis: Recent theoretical progress has established local PL conditions for two-layer linear networks under general losses and relaxed assumptions on step size, width, and initialization (Xu et al., 2025).
+Moreover, mild over-parameterization alone suffices for linear convergence in two-layer nets, as shown by (Zhou et al., 2021).
 However, this analysis is limited to linear networks and does not extend to the non-linear case that dominates practical applications.
@@
 2.2 Neural Tangent Kernel Theory and Finite-Width Analysis
-The Neural Tangent Kernel framework, introduced by Jacot2018 , revolutionized the theoretical understanding of neural network training by showing that infinitely wide networks behave as kernel methods.
-This insight has been crucial for proving global convergence guarantees, but typically requires unrealistic infinite-width assumptions Jacot2018 ; novak2022fast .
+The Neural Tangent Kernel framework, introduced by (Jacot et al., 2018), revolutionized the theoretical understanding of neural network training by showing that infinitely wide networks behave as kernel methods.
+This insight has been crucial for proving global convergence guarantees, but typically requires unrealistic infinite-width assumptions (Jacot et al., 2018; Novak et al., 2022).
 Finite-Width Extensions: Recent work has sought to extend NTK insights to finite-width settings.
-AllenZhu2019 showed that sufficiently wide networks can maintain NTK stability, while Arora2019a characterized when finite networks approximate the infinite-width behavior Arora2019a ; AllenZhu2019 ; frei2022proxy .
+(Allen-Zhu et al., 2019) showed that sufficiently wide networks can maintain NTK stability, while (Arora et al., 2019) characterized when finite networks approximate the infinite-width behavior (Arora et al., 2019; Allen-Zhu et al., 2019; Frei and Gu, 2022).
 However, these results often require exponentially large widths to provide meaningful guarantees.
@@
 Understanding the geometric structure of neural network loss landscapes has been crucial for explaining optimization success.
-choromanska2015loss showed that poor local minima are exponentially rare using spin-glass models, while kawaguchi2016deep proved that all local minima are global for deep linear networks.
+(Choromanska et al., 2015) showed that poor local minima are exponentially rare using spin-glass models, while (Kawaguchi, 2016) proved that all local minima are global for deep linear networks.
 Local Geometric Analysis: Recent work has developed general mathematical frameworks for loss landscapes in over-parameterized systems, providing modern views of efficient optimization.
-The LQCR framework Aich2025 contributes to this line by providing concrete geometric conditions that ensure both convergence and fast rates.
+The LQCR framework (Aich et al., 2025) contributes to this line by providing concrete geometric conditions that ensure both convergence and fast rates.
 Connection to Practice: A key challenge has been connecting geometric insights to practical training dynamics.
 Our work addresses this by showing that the same geometric properties (LQCR structure) that ensure convergence also provide the functional properties (local PL condition) needed for fast optimization.
-3.  Notation
+3 Notation
 In this section, we introduce the notation used throughout our analysis.
-We consider an optimization problem for deep neural networks with the following symbols: $f_{\theta}$, $\theta$, $p$, $D$, $m$, $x$, $d$, $y$, $n$, $\mathcal{L}$, $\ell$, $\mathcal{R}$, $\eta$, $t$, $L$, $L_{\Theta}$, $\alpha$, $\gamma$, $\mu$, $\mathcal{L}^{*}$, $\mathcal{L}_{\mathcal{R}}^{*}$, $J_{\theta}$, $\mathbf{\Theta}_{\theta}$, and $\lambda_{\min}$.
-[…]
-4.  From Local Geometry to the Polyak-Lojasiewicz Condition
+We consider an optimization problem for deep neural networks with the following symbols: $f_{\theta}$, $\theta$, $p$, $D$, $m$, $x$, $d$, $y$, $n$, $\mathcal{L}$, $\ell$, $\mathcal{R}$, $\eta$, $t$, $L$, $L_{\Theta}$, $\alpha$, $\gamma$, $\mu$, $\mathcal{L}^{*}$, $\mathcal{L}_{\mathcal{R}}^{*}$, $J_{\theta}$, $\boldsymbol{\theta}_{\theta}$, and $\lambda_{\min}$.
+Let $f_{\theta}$ be a deep neural network with $D$ layers, width $m$, and parameters $\theta\in\mathbb{R}^{p}$.
+For a dataset with $n$ samples of input-label pairs $(x,y)$, where $x\in\mathbb{R}^{d}$, we aim to minimize the empirical loss function $\mathcal{L}(\theta)$, which is the average of a per-sample loss $\ell$.
+The gradient descent algorithm proceeds over iterations $t$ with a learning rate $\eta$.
+Our analysis takes place within a local region near initialization, denoted by $\mathcal{R}\subset\mathbb{R}^{p}$.
+Our theory connects several key concepts.
+The properties of a Locally Quasi-Convex Region (LQCR) are defined by a curvature parameter $\alpha$ and a descent parameter $\gamma$.
+The Polyak-Lojasiewicz (PL) condition is defined by a constant $\mu$, the global minimum loss value $\mathcal{L}^{*}$, and the regional minimum loss $\mathcal{L}_{\mathcal{R}}^{*}$.
+The loss function $\mathcal{L}$ is assumed to be $L$-smooth.
+Finally, our link to deep learning theory relies on the network’s Jacobian $J_{\theta}$, the resulting Neural Tangent Kernel (NTK) matrix $\boldsymbol{\theta}_{\theta}$ (which is assumed to be $L_{\Theta}$-Lipschitz), and its smallest eigenvalue, which we assume is bounded below by $\lambda_{\min}>0$ within the region $\mathcal{R}$.
+4 From Local Geometry to the Polyak-Lojasiewicz Condition
 In this section, we introduce the concepts required to link the local geometry of a neural network’s loss landscape to a condition sufficient for linear convergence.
@@
 One can tackle this problem by demonstrating that the well-behaved local regions established in prior work also satisfy the Polyak-Lojasiewicz (PL) condition locally, a property we formalize below.
-The PL condition, first introduced by Polyak Polyak1963 , is a powerful tool for proving linear convergence on non-convex functions. However, the global PL condition, defined as
-$\frac{1}{2}||\nabla\mathcal{L}(\theta)||^{2}\geq\mu(\mathcal{L}(\theta)-\mathcal{L}^{*})$ (1)
+The PL condition, first introduced by Polyak (Polyak, 1963), is a powerful tool for proving linear convergence on non-convex functions. However, the global PL condition, defined as
+[…]
 where $\mathcal{L}^{*}$ is the global minimum, is too restrictive for most DNNs. This motivates a localized version that only needs to hold within the region of optimization.
-Definition 4.1 (Locally Polyak-Lojasiewicz (LPL) Region)
+Definition 4.1 (Locally Polyak-Lojasiewicz (LPL) Region).
 A region $\mathcal{R}\subset\mathbb{R}^{p}$ is a $(\mu,\mathcal{R})$-Locally Polyak-Lojasiewicz Region (LPLR) if the PL condition holds for all $\theta\in\mathcal{R}$:
-$\frac{1}{2}||\nabla\mathcal{L}(\theta)||^{2}\geq\mu(\mathcal{L}(\theta)-\mathcal{L}_{\mathcal{R}}^{*})\quad\forall\theta\in\mathcal{R}$ (2)
+[…]
 where $\mathcal{L}_{\mathcal{R}}^{*}=\min_{\theta\in\mathcal{R}}\mathcal{L}(\theta)$ is the minimum value of the loss within the region $\mathcal{R}$.
 The existence of such an LPLR is not arbitrary; it is a direct consequence of the underlying geometry of the region $\mathcal{R}$.
-Specifically, this analysis builds upon the concept of a Locally Quasi-Convex Region (LQCR) introduced in Aich2025 .
+Specifically, this analysis builds upon the concept of a Locally Quasi-Convex Region (LQCR) introduced in (Aich et al., 2025).
 An LQCR is a region with bounded negative curvature and a guaranteed descent direction, properties that ensure iterates make steady progress.
-Definition 4.2 (Locally Quasi-Convex Region Aich2025 )
+Definition 4.2 (Locally Quasi-Convex Region (Aich et al., 2025)).
 A region $\mathcal{R}\subset\mathbb{R}^{p}$ is an $(\alpha,\gamma)$-LQCR if for all $\theta,\phi\in\mathcal{R}$:
-  1. i.
-    Curvature Bound: $\mathcal{L}(\phi)\geq\mathcal{L}(\theta)+\nabla\mathcal{L}(\theta)^{\top}(\phi-\theta)-\frac{\alpha}{2}||\phi-\theta||^{2}$.
-  2. ii.
-    Descent Condition: If $\mathcal{L}(\phi)<\mathcal{L}(\theta)$, then $\langle\nabla\mathcal{L}(\theta),\phi-\theta\rangle\geq\gamma||\phi-\theta||||\nabla\mathcal{L}(\theta)||$.
-The critical link between the geometry of an LQCR and the functional property of an LPLR is the Neural Tangent Kernel (NTK) Jacot2018 .
+  1. 1.
+[…]
+  2. 2.
+[…]
+The critical link between the geometry of an LQCR and the functional property of an LPLR is the Neural Tangent Kernel (NTK) (Jacot et al., 2018).
 The empirical NTK, $\Theta_{\theta}$, is defined from the network Jacobian $J_{\theta}(x)$ as $\Theta_{\theta}(x,x^{\prime})=J_{\theta}(x)J_{\theta}(x^{\prime})^{\top}$.
 For our analysis to hold, we assume that this kernel remains well-behaved within the LQCR.
-This local stability is expected in many modern deep learning scenarios, including sufficiently overparameterized or wide networks, and when using standard initialization schemes and constrained learning rates Arora2019a ; AllenZhu2019 ; Aich2025 .
+This local stability is expected in many modern deep learning scenarios, including sufficiently overparameterized or wide networks, and when using standard initialization schemes and constrained learning rates (Arora et al., 2019; Allen-Zhu et al., 2019; Aich et al., 2025).
 One can show that an LQCR satisfying this local NTK stability is also an LPLR.
@@
 Specifically, for a point $\theta\in\mathcal{R}$ where the smallest eigenvalue of the NTK, $\lambda_{\min}(\Theta_{\theta})$, is positive, the following inequality holds:
-$\mathcal{L}(\theta)-\mathcal{L}_{\mathcal{R}}^{*}\leq\frac{1}{2\lambda_{\min}(\Theta_{\theta})}||\nabla\mathcal{L}(\theta)||^{2}$ (3)
+[…]
 This relationship directly gives rise to the local PL condition.
 By rearranging the inequality, we find that the PL constant $\mu$ is precisely the smallest eigenvalue of the NTK, $\lambda_{\min}$, within the stable region $\mathcal{R}$.
-5.  Main Theoretical Results
+Remark 4.3 (On the descent condition orientation).
+In Definition 4.2, the descent condition is stated using the displacement vector $(\phi-\theta)$.
+This choice is without loss of generality: equivalently, one could write $(\theta-\phi)$ with the inequality sign flipped.
+The essential requirement is that, whenever $\mathcal{L}(\phi)<\mathcal{L}(\theta)$, the gradient at $\theta$ has a positive projection along the displacement toward $\phi$.
+This convention ensures the geometry is consistent and avoids confusion about sign.
+For consistency with this convention, we restate the descent condition from prior LQCR work (Aich et al., 2025); the equivalent form depends on whether the displacement is taken as $(\phi-\theta)$ or $(\theta-\phi)$.
+5 Main Theoretical Results
 In this section, we present our main theorems, which provide formal guarantees for the existence of Locally Polyak-Lojasiewicz Regions (LPLRs) and prove the resulting linear convergence of gradient descent.
@@
 Our central result establishes that an LQCR is also an LPLR.
-Theorem 5.1 (Existence of LPLRs)
-Under the same setup as prior work on LQCRs Aich2025 and holding Assumption 5.2, the LQCR $\mathcal{R}$ is also a Local Polyak-Lojasiewicz Region (LPLR).
+Theorem 5.1 (Existence of LPLRs).
+Under the same setup as prior work on LQCRs (Aich et al., 2025) and holding Assumption 5.2, the LQCR $\mathcal{R}$ is also a Local Polyak-Lojasiewicz Region (LPLR).
 Specifically, the loss $\mathcal{L}$ satisfies the PL condition within $\mathcal{R}$ for a PL constant $\mu$ satisfying:
-$\mu\geq\lambda_{\min}$ (4)
+$\mu=\lambda_{\min}$ (4)
 where $\lambda_{\min}$ is the lower bound on the smallest eigenvalue of the NTK from Assumption 5.2.
@@
 We state this formally below.
-Assumption 5.2 (Enhanced Local NTK Stability)
-Let $\mathcal{R}$ be the $(\alpha,\gamma)$-LQCR established in Aich2025 . We assume that for all $\theta\in\mathcal{R}$, the empirical Neural Tangent Kernel $\mathbf{\Theta}_{\theta}$ satisfies:
-  • Positive Definiteness: $\mathbf{\Theta}_{\theta}\succ 0$ with smallest eigenvalue $\lambda_{\min}(\mathbf{\Theta}_{\theta})\geq\lambda_{\min}>0$.
+Assumption 5.2 (Enhanced Local NTK Stability).
+Let $\mathcal{R}$ be the $(\alpha,\gamma)$-LQCR established in (Aich et al., 2025).
+We assume that for all $\theta\in\mathcal{R}$, the empirical Neural Tangent Kernel $\boldsymbol{\theta}_{\theta}$ satisfies:
+  • Positive Definiteness: $\boldsymbol{\theta}_{\theta}\succ 0$ with smallest eigenvalue $\lambda_{\min}(\boldsymbol{\theta}_{\theta})\geq\lambda_{\min}>0$.
   • Stability Condition: For any $\theta,\phi\in\mathcal{R}$:
-    $\|\mathbf{\Theta}_{\theta}-\mathbf{\Theta}_{\phi}\|\leq L_{\Theta}\|\theta-\phi\|$ (5)
+    $\|\boldsymbol{\theta}_{\theta}-\boldsymbol{\theta}_{\phi}\|\leq L_{\Theta}\|\theta-\phi\|$ (5)
     where $L_{\Theta}$ is the NTK Lipschitz constant within $\mathcal{R}$.
-Sufficient Conditions: This assumption holds when the network width $m$, initialization, learning rate $\eta$, and region radius $R_{\mathcal{R}}$ satisfy certain criteria, such as those outlined in recent theoretical work on over-parameterized models.
+Sufficient Conditions: This assumption holds when the network width $m$, initialization, learning rate $\eta$, and region radius $r(R)$ satisfy certain criteria, such as those outlined in recent theoretical work on over-parameterized models.
 Proof of Theorem 5.1.
@@
 The squared norm of the gradient is $\|\nabla\mathcal{L}(\theta)\|^{2}=\frac{1}{n^{2}}\|J_{\theta}^{T}(f_{\theta}-y)\|^{2}$.
-The empirical Neural Tangent Kernel (NTK) is defined as $\mathbf{\Theta}_{\theta}=\frac{1}{n}J_{\theta}J_{\theta}^{T}$.
+The empirical Neural Tangent Kernel (NTK) is defined as $\boldsymbol{\theta}_{\theta}=\frac{1}{n}J_{\theta}J_{\theta}^{T}$.
 The smallest non-zero eigenvalue of $J_{\theta}^{T}J_{\theta}$ is equal to that of $J_{\theta}J_{\theta}^{T}$.
-This property, combined with Assumption 5.2 which guarantees $\lambda_{\min}(\mathbf{\Theta}_{\theta})\geq\lambda_{\min}>0$, gives us a bound on the squared gradient:
-$\|J_{\theta}^{T}(f_{\theta}-y)\|^{2}\geq\lambda_{\min}(J_{\theta}^{T}J_{\theta})\|f_{\theta}-y\|^{2}=n\lambda_{\min}(\mathbf{\Theta}_{\theta})\|f_{\theta}-y\|^{2}\geq n\lambda_{\min}\|f_{\theta}-y\|^{2}.$ (6)
+This property, combined with Assumption 5.2 which guarantees $\lambda_{\min}(\boldsymbol{\theta}_{\theta})\geq\lambda_{\min}>0$, gives us a bound on the squared gradient:
+$\|J_{\theta}^{T}(f_{\theta}-y)\|^{2}\geq\lambda_{\min}(J_{\theta}^{T}J_{\theta})\|f_{\theta}-y\|^{2}$ (6)
+$=n\lambda_{\min}(\boldsymbol{\theta}_{\theta})\|f_{\theta}-y\|^{2}\geq n\lambda_{\min}\|f_{\theta}-y\|^{2}.$ (7)
 Substituting this back into the expression for the gradient norm, we get:
@@
 $\frac{1}{2}\|\nabla\mathcal{L}(\theta)\|^{2}\geq\lambda_{\min}(\mathcal{L}(\theta)-\mathcal{L}_{\mathcal{R}}^{*}).$ (10)
-This establishes the result with a local PL constant $\mu=\lambda_{\min}$.
-Remark 5.3
+This establishes the result with a local PL constant $\mu=\lambda_{\min}$. ∎
+Remark 5.3.
 While global NTK stability requires unrealistic infinite-width networks, our assumption is substantially weaker, requiring only that the kernel remains well-conditioned within the local optimization region $\mathcal{R}$.
-This aligns with the "lazy training" phenomenon, where parameters in sufficiently overparameterized networks stay close to initialization Jacot2018 .
+This aligns with the ”lazy training” phenomenon, where parameters in sufficiently overparameterized networks stay close to initialization (Jacot et al., 2018).
 Our contribution is to show that this empirically observed local stability is the critical ingredient that transitions the convergence rate from sublinear to linear.
@@
 This property makes the LPLR a powerful tool for explaining the efficiency of deep learning optimizers.
-Theorem 5.4 (Linear Convergence of Gradient Descent)
+Theorem 5.4 (Linear Convergence of Gradient Descent).
 Assume the loss function $\mathcal{L}$ is L-smooth and the conditions of Theorem 5.1 hold.
@@
 Proof of Theorem 5.4.
-The proof is a direct application of the standard result for gradient descent on a function satisfying the PL condition, as shown by Polyak in 1963 Polyak1963 and reviewed in 2016 by Karimi et al.Karimi2016 .
+The proof is a direct application of the standard result for gradient descent on a function satisfying the PL condition, as shown by Polyak in 1963 (Polyak, 1963) and reviewed in 2016 by Karimi et al.(Karimi et al., 2016).
 The standard descent lemma for an L-smooth function states that for a step size $\eta=1/L$:
-$\mathcal{L}(\theta^{(t+1)})\leq\mathcal{L}(\theta^{(t)})-\frac{1}{2L}||\nabla\mathcal{L}(\theta^{(t)})||^{2}.$ (11)
+[…]
 From Theorem 5.1, we know the local PL condition holds with constant $\mu=\lambda_{\min}$:
-$||\nabla\mathcal{L}(\theta^{(t)})||^{2}\geq 2\lambda_{\min}(\mathcal{L}(\theta^{(t)})-\mathcal{L}_{\mathcal{R}}^{*}).$ (12)
+[…]
 Substituting (13) into (12) gives:
@@
 $\mathcal{L}(\theta^{(t+1)})-\mathcal{L}_{\mathcal{R}}^{*}\leq(1-\frac{\lambda_{\min}}{L})(\mathcal{L}(\theta^{(t)})-\mathcal{L}_{\mathcal{R}}^{*}).$ (15)
-Applying this inequality recursively from $t=0$ yields the final result. For a general step size $\eta\leq 1/L$, the rate becomes $(1-\eta\lambda_{\min})^{t}$.
-6.  Empirical Validation
+Applying this inequality recursively from $t=0$ yields the final result. For a general step size $\eta\leq 1/L$, the rate becomes $(1-\eta\lambda_{\min})^{t}$. ∎
+6 Empirical Validation
 In this section, we present a series of experiments designed to empirically validate our theoretical findings, confirming the existence of LPLRs and the resulting linear convergence in both controlled and realistic settings.
-The aim is to first verify our core claims in a controlled setting using a fully-connected network with full-batch Gradient Descent, and then to test the robustness of our theory in a more complex scenario involving a ResNet-style CNN trained with stochastic methods.
+Our analysis focuses on the transition from initial optimization dynamics into the stable, fast-convergence regime predicted by our theory.
 6.1 Validation of the Linear Convergence Rate
-Our primary validation uses a 5-layer fully-connected network with a width of 2048 and ReLU activations, trained on a binary subset of the MNIST dataset Nair2010 ; He2015 ; LeCun1998 .
-To precisely match our theoretical setting, we use full-batch Gradient Descent with a learning rate of 0.001 Polyak1963 ; Nesterov2004 .
-The regional minimum loss, $\mathcal{L}_{\mathcal{R}}^{*}$, is estimated by taking the minimum training loss observed over the full 250 epochs.
-While this is an empirical proxy for the true regional minimum, it is a standard practice in convergence analysis Nesterov2004 ; Johnson2013 and provides a close and practical estimate for calculating the suboptimality gap.
-In this controlled environment, our experiments perfectly align with our theory.
-As shown in Figure 1, the suboptimality gap decays linearly on a semi-log plot, where the straight line provides clear evidence of the linear convergence rate predicted by Theorem 5.4.
-This result confirms that the LPLR structure leads to fast optimization.
+Our primary validation uses a 5-layer fully-connected network with a width of 2048 and ReLU activations, trained on a binary subset of the MNIST dataset (Nair and Hinton, 2010; He et al., 2015; LeCun et al., 1998).
+To match our theoretical setting, we use full-batch Gradient Descent with a learning rate of 0.001 (Polyak, 1963; Nesterov, 2004).
+The regional minimum loss, $\mathcal{L}_{\mathcal{R}}^{*}$, is estimated using the final training loss at the end of the 250-epoch run.
+This provides a consistent empirical proxy for the terminal loss level used to compute the suboptimality gap.
+As shown in Figure 1, the suboptimality gap decreases steadily on a semi-log scale over the course of training, with an initial transient followed by a long regime of rapid decay.
+While the curve is not perfectly straight over the entire trajectory (reflecting changing local contraction behavior and the use of a terminal-loss proxy for $\mathcal{L}_{\mathcal{R}}^{*}$), the overall dynamics are consistent with a fast, near-linear convergence phase as predicted by Theorem 5.4.
 Figure 1: Training loss dynamics for the MLP on MNIST.
 The y-axis (suboptimality gap) is on a logarithmic scale.
-The observed straight line provides strong evidence for the linear convergence rate predicted by Theorem 5.4.
+After an initial transient, the gap decays rapidly over a long training regime, consistent with the linear-rate behavior predicted by Theorem 5.4.
+6.2 Comparative Analysis of Initialization Strategies
+Next, we compare standard He initialization to an enhanced, depth-aware initialization strategy1111The enhanced initialization, proposed by (Aich et al., 2025), scales the variance of the standard He initialization for each layer $l$ by a factor $g(l,D)=(1-|D/2-l|/D)^{2}$, where $D$ is the total network depth..
+Figure 2 shows that both initializations enter a similar rapid-decay regime over training.
+A visible difference occurs early in optimization: the standard He initialization exhibits a sharper initial transient, whereas the enhanced initialization yields a smoother start.
+Beyond the initial phase, the two trajectories remain very close, indicating that (for this setup) the long-run convergence behavior is broadly robust to these two initializations.
+Figure 2: Comparison of convergence dynamics for standard and enhanced initializations for the MLP on MNIST (semi-log scale).
+The enhanced initialization reduces the severity of the early transient, while both methods display very similar decay behavior across most of training.
 6.3 Verification of the Local PL Condition
-To further investigate the mechanism driving this convergence, we test for the local PL-like property.
-The results, shown in Figure 2, verify its existence.
-The observed linear relationship on the log-log plot (slope 0.71$\approx$ 0.71≈ 0.71) confirms that the squared gradient norm is polynomially bounded by the suboptimality gap.
-This provides strong evidence for a local Kurdyka-Łojasiewicz (KL) property, which is sufficient to drive linear convergence Karimi2016 .This provides strong evidence for a local Kurdyka-Łojasiewicz (KL) property, which is sufficient to drive linear convergence.
-While the theoretical PL condition corresponds to a slope of 1, slopes less than 1 are common in practice and still indicate a strong polynomial relationship that guarantees fast convergence.
-Figure 2: Verification of the local PL condition for the MLP on MNIST.
-The linear relationship on the log-log scale confirms that $||\nabla\mathcal{L}(\theta)||^{2}$ is lower-bounded by a multiple of $(\mathcal{L}(\theta)-\mathcal{L}_{\mathcal{R}}^{*})$, a behavior consistent with PL-like conditions known to guarantee linear convergence Necoara2015 ; Richtarik2014 .
+To probe the mechanism behind the fast decay observed above, we examine the empirical relationship between the squared gradient norm and the suboptimality gap.
+Figure 3 presents a log-log plot of $\nabla\mathcal{L}(\theta)||^{2}$ against $(\mathcal{L}(\theta)-\mathcal{L}_{\mathcal{R}}^{*})$ (with $\mathcal{L}_{\mathcal{R}}^{*}$ again taken as the final training loss).
+The fitted trend line in the figure has slope approximately $1.14$, indicating a strong polynomial coupling between gradient magnitude and suboptimality in the explored region.
+This behavior is consistent with PL/KL-type error-bound phenomena associated with fast convergence (Karimi et al., 2016), and supports the view that optimization is proceeding in a locally favorable region where gradients remain informative about function suboptimality.
+Figure 3: Empirical relationship between squared gradient norm and suboptimality gap for the MLP on MNIST (log-log scale).
+The strong polynomial coupling (with fitted slope $\approx 1.14$ over the plotted range) is consistent with PL/KL-type error-bound behavior known to yield fast rates (Necoara et al., 2019; Richtárik and Takáč, 2014).
 6.4 Generalization to Complex Architectures
-To demonstrate the robustness and generality of our theory, we conduct a second experiment in a more challenging setting.
-We train a ResNet-style CNN on a 5-class subset of the CIFAR-10 dataset Krizhevsky2009 .
-We use mini-batch Stochastic Gradient Descent (SGD) with a batch size of 128 and a cosine annealing learning rate schedule, which are standard practices for training modern CNNs.
-Even with the added complexity of a deeper architecture and the noise from stochastic optimization, the core predictions of our theory hold.
-Figure 3 (left) shows the smoothed training loss, which again exhibits a clear linear decay on the semi-log plot.
-The right panel confirms that a PL-like property emerges and drives fast convergence even in this more complex setting, with the smoothed data revealing a strong linear trend (slope 0.77$\approx$ 0.77≈ 0.77).
-Figure 3: Validation on a ResNet-style CNN trained on a CIFAR-10 subset with SGD.
-(Left) The smoothed training loss shows clear linear convergence on a semi-log plot.
-(Right) The smoothed PL-condition plot confirms a strong linear relationship (slope 0.77$\approx$ 0.77≈ 0.77), demonstrating that our theory generalizes to complex architectures and stochastic optimization.
-6.4 Comparative Analysis of Initialization Strategies
-Finally, we provide a comparative analysis of the standard He initialization versus an enhanced, depth-aware initialization strategy [^1] The enhanced initialization, proposed by Aich2025 , scales the variance of the standard He initialization for each layer $l$ by a factor $g(l,D)=(1-|D/2-l|/D)^{2}$, where $D$ is the total network depth.
-The goal of this method is to balance signal propagation..
-As shown in Figure 4, while both methods exhibit linear convergence, the enhanced method, which scales down the initial variance, resulted in a slower convergence rate.
-This suggests that for this architecture, the larger initial weights provided by the standard He initialization create a more favorable landscape for rapid optimization, highlighting that the ideal initialization strategy is highly dependent on the specific problem setting.
-Figure 4: Comparison of convergence rates for standard and enhanced initializations for the MLP on MNIST.
-While both methods exhibit linear convergence, the standard He initialization converges significantly faster.
-This suggests that reducing the initial weight variance, as done by the enhanced method, can impede the speed of optimization for this architecture.
+To demonstrate the robustness and generality of our theory beyond the controlled full-batch setting, we repeat the empirical validation on a substantially more challenging training pipeline.
+We train a ResNet-style CNN on a 5-class subset of the CIFAR-10 dataset (Krizhevsky and Hinton, 2009) using mini-batch Stochastic Gradient Descent (SGD) with batch size 128 and a cosine annealing learning-rate schedule, which are standard choices for modern convolutional architectures.
+Despite the deeper architecture and the additional noise induced by stochastic optimization, qualitative signatures consistent with our theory remain visible.
+Figure 4 (left) shows the smoothed suboptimality gap on a semi-log scale, exhibiting an extended regime of approximately linear decay, which is consistent with linear-rate behavior over the explored region.
+The right panel provides complementary evidence of an error-bound mechanism emerging in this setting: the log-log relationship between the squared gradient norm and the suboptimality gap is well-approximated by a power-law trend over a substantial range, with an empirical fitted slope of approximately $0.29$.
+While this value is below the ideal PL scaling of $1$, it nevertheless indicates a nontrivial polynomial coupling between gradient magnitude and suboptimality, consistent with the presence of a PL/KL-like error-bound effect in this more complex, stochastic training pipeline.
+Figure 4: Validation on a ResNet-style CNN trained on a 5-class CIFAR-10 subset with SGD (smoothed curves).
+(Left) The smoothed suboptimality gap decays approximately linearly on a semi-log scale over an extended regime, indicating linear-rate behavior in the explored region.
+(Right) The log-log plot of squared gradient norm versus suboptimality gap exhibits an approximately power-law trend; the fitted slope is $\approx 0.29$, providing evidence of a PL/KL-like error-bound relationship in this complex, stochastic setting.
 6.5 Ablation Study: Impact of Network Width
-Our theory establishes that sufficiently over-parameterized networks exhibit local regions of favorable geometry (LPLRs), leading to linear convergence.
-A key parameter governing over-parameterization is network width.
-To empirically investigate its impact in our setting, we conducted an ablation study.
-[…]
-Figure 5: Impact of network width on the final loss value achieved by gradient descent after 250 epochs.
-The result demonstrates that wider networks consistently find solutions with lower training loss.
-Results and Discussion: As shown in Figure 5, there is a clear and consistent trend: as network width increases, the final training loss achieved by the optimizer decreases.
-This finding aligns with established theories of over-parameterization, which suggest that wider models have smoother, more favorable loss landscapes.
-While our main theorems guarantee a linear convergence rate once inside an LPLR, this ablation study suggests that wider models may converge to "better" LPLRs, i.e., those surrounding a lower-loss minimum.
-This provides an important empirical link between the global effects of over-parameterization and the local convergence dynamics central to our work.
-7.  Discussion and Conclusion
-This paper provides a definitive answer to one of the most fundamental questions in deep learning theory: why do gradient-based methods converge so rapidly in practice, despite the highly non-convex nature of neural network loss landscapes?
-Our work closes the critical gap between the sublinear convergence rates guaranteed by prior geometric analyses and the exponential rates consistently observed in practice.
-We accomplish this by demonstrating that the Locally Quasi-Convex Regions (LQCRs) Aich2025 that ensure convergence also exhibit a local Polyak-Lojasiewicz property Polyak1963 under a mild and realistic assumption about Neural Tangent Kernel stability Jacot2018 .
-7.1 Comparison with Recent Advances
-Architectural Approaches: Recent work has shown that networks can be designed to satisfy PL conditions by construction (PLNet) wang2024monotone .
-While this guarantees fast convergence, it requires abandoning standard architectures.
-Our approach explains why standard architectures already exhibit these favorable properties.
-Stochastic Analysis: an2024convergence provided convergence analysis for SGD under local Lojasiewicz conditions, but their analysis requires stronger assumptions about the local landscape.
-Our geometric approach provides a path toward understanding when these conditions hold.
-Linear Network Insights: Recent progress on linear networks xu2025local shows that local PL properties can be established under relaxed assumptions.
-Our work extends these insights to the non-linear case that dominates practical applications.
-7.2 Theoretical Contributions and Significance
-Our central theoretical contribution is the introduction and rigorous analysis of the Locally Polyak-Lojasiewicz Region (LPLR).
-This concept represents a significant advance in understanding neural network optimization by showing that:
-  • Local geometry implies fast optimization: The same well-behaved regions that guarantee monotonic progress also satisfy a condition sufficient for linear convergence.
-  • Finite-width networks can achieve linear rates: Unlike prior work requiring infinite width, our theory applies to practically-sized networks under realistic conditions.
-  • The NTK provides the crucial link: Local stability of the Neural Tangent Kernel, a much weaker condition than global stability, is sufficient to induce the PL property that drives fast convergence.
-Our theoretical framework elegantly unifies three previously disparate lines of research in deep learning theory:
-  • Local Landscape Analysis (LQCRs): Provides the structural guarantee that optimization remains within a well-behaved region.
-  • Neural Tangent Kernel Theory: Contributes the mechanism (local kernel stability) that endows this region with favorable optimization properties.
-  • Fast Optimization Theory (PL Condition): Supplies the analytical tools to translate geometric properties into convergence rate guarantees.
-This synthesis represents a qualitative advance in our theoretical understanding, moving beyond existence proofs to provide concrete, quantitative explanations for the efficiency of deep learning optimization.
-7.3 Empirical Validation and Practical Relevance
-Our empirical validation demonstrates that the LPLR framework is not merely a theoretical construct but captures genuine phenomena occurring in practical deep learning.
-We validated our theory across multiple settings:
-  • Controlled validation: Full-batch gradient descent on fully-connected networks provided exact confirmation of our theoretical predictions.
-  • Realistic scenarios: ResNet-style CNNs trained with mini-batch SGD on CIFAR-10 showed that the same principles govern modern architectures and stochastic optimization.
-  • Initialization sensitivity: A comparison of initialization schemes highlighted the critical impact of initial weight variance on convergence speed, demonstrating the practical importance of landscape geometry at the start of training.
-The robustness of our findings across these diverse settings strongly suggests that LPLRs capture a fundamental organizing principle of neural network optimization, rather than an artifact of specific architectures or training procedures.
-7.4 Implications for Deep Learning Practice
-Our work has several important implications for practitioners:
-  • Initialization Design: Our theory provides a principled framework for analyzing initialization schemes.
-The goal is to find methods that improve the local PL constant $\mu$, as this should translate to faster training.
-However, as our comparative experiment shows, the relationship between a specific strategy and the resulting landscape geometry is complex and architecture-dependent, making this a critical area for future investigation.
-  • Architecture Design: Understanding that local NTK stability drives fast convergence can inform architectural choices.
-Designs that promote better conditioning of the local kernel (through appropriate width, normalization, or connectivity patterns) should exhibit superior optimization properties.
-  • Learning Rate Selection: Our linear convergence guarantee provides theoretical backing for aggressive learning rate schedules, as long as they respect the smoothness constraints and maintain iterates within the LPLR.
-  • Optimization Algorithm Development: The local PL property suggests that algorithms specifically designed to maintain or enhance this property could achieve superior performance.
-7.5 Limitations and Scope
-While our work represents a significant advance, we acknowledge several limitations that define important directions for future research:
-  • Stochastic Analysis: Our main theoretical results are established for full-batch gradient descent.
-Although our empirical results strongly suggest the theory extends to stochastic settings, a formal analysis of SGD convergence to a noise ball within an LPLR remains an important open problem.
-  • NTK Assumption: Our theory relies on local NTK stability, which we validate empirically but do not fully characterize theoretically.
-Understanding precisely when this assumption holds across different architectures and training regimes would strengthen our framework.
-  • Generalization Connection: While our work explains training efficiency, the connection between LPLR properties and final generalization performance remains unexplored.
-This represents a promising avenue for understanding implicit regularization.
-7.6 Future Research Directions
-Our work opens several exciting research directions:
-  • Stochastic Extension: Developing a rigorous theory for mini-batch SGD within LPLRs, characterizing both the convergence rate and the size of the noise ball, building on prior analyses of stochastic methods under PL-like conditions Karimi2016 ; Kingma2015 ; Reddi2018 , would significantly extend the practical relevance of our framework.
-  • Architectural Analysis: Investigating how specific architectural components (attention mechanisms, normalization layers, skip connections) affect the PL constant $\mu$ could provide concrete design principles for faster-training networks.
-  • Adaptive Optimization: Developing optimization algorithms that dynamically estimate and exploit the local PL constant could lead to more efficient training procedures.
-  • Generalization Theory: Connecting LPLR properties to generalization bounds could provide new insights into why certain training trajectories lead to better test performance.
-  • Empirical Scaling Laws: Characterizing how the size and properties of LPLRs scale with network width, depth, and dataset size could inform resource allocation decisions in large-scale training.
-7.7 Broader Impact
-Our work contributes to the theoretical foundations of deep learning by providing a principled explanation for one of its most important empirical phenomena.
-By showing that local geometric properties naturally give rise to favorable optimization conditions, we advance the field’s understanding of why deep learning works so well in practice.
-This theoretical clarity can guide both algorithm development and architectural design, potentially leading to more efficient training procedures and better-performing models.
-Furthermore, by demonstrating that finite-width networks can achieve the same favorable convergence guarantees as their infinite-width counterparts under realistic conditions, our work helps bridge the gap between theoretical analysis and practical implementation.
-This has important implications for the design of resource-efficient training procedures and for understanding the trade-offs between model size and optimization efficiency.
-7.8 Conclusion
-This paper resolves a longstanding puzzle in deep learning theory by showing that the rapid convergence observed in practice emerges naturally from the local geometric structure of neural network loss landscapes.
-Our introduction of Locally Polyak-Lojasiewicz Regions provides both a theoretical framework for understanding this phenomenon and practical insights for improving optimization in deep learning.
-The convergence of geometric, kernel-based, and optimization-theoretic perspectives in our analysis represents a significant step toward a more complete understanding of why deep learning optimization is so remarkably effective.
-The validation of our theory across both controlled and realistic settings, combined with its connection to fundamental concepts in optimization and deep learning theory, positions this work as a significant contribution to the field.
-As deep learning continues to scale to ever-larger models and datasets, theoretical frameworks like ours that explain and predict the behavior of optimization algorithms become increasingly valuable for both understanding current successes and guiding future developments.
-Data and Code Availability
-The MNIST and CIFAR-10 datasets used in this study are publicly available through the torchvision library.
-The code to reproduce all experiments and figures presented in this paper will be made available on GitHub upon publication.
+Our theory predicts that sufficient over-parameterization promotes favorable local geometry (LPLRs), which in turn supports fast optimization dynamics.
+Since width is a primary driver of over-parameterization in fully-connected networks, we conduct a controlled ablation study to quantify how width affects the optimization outcome in our experimental setting.
+Experimental Setup: We use the same full-batch Gradient Descent configuration as in our main MNIST experiments: a 5-layer ReLU MLP trained on a binary MNIST classification task with squared loss.
+We vary the hidden-layer width over $\{512,1024,2048,4096\}$, while keeping all other hyperparameters fixed (depth, learning rate, and number of training epochs).
+For each width, we train for 250 epochs and record the final training loss.
+Table 1: Width ablation results for the 5-layer MLP on binary MNIST (full-batch GD, 250 epochs). Wider networks achieve lower final training loss under identical optimization settings.
+Width Final training loss
+512 0.054571
+1024 0.045054
+2048 0.037100
+4096 0.031133
+Figure 5: Impact of network width on the final training loss achieved after 250 epochs. The trend is monotone: larger widths consistently yield lower final loss.
+Results and Discussion: The results show a clear and consistent monotone improvement as width increases.
+Table 1 reports the exact final losses, and Figure 5 visualizes the same trend.
+Under identical training conditions, increasing width from 512 to 4096 reduces the final loss from 0.054571 to 0.031133.
+This empirical pattern supports the intuition that greater over-parameterization makes it easier for gradient-based optimization to reach lower-loss solutions in practice.
+While our main theorems address the rate of convergence once the iterates enter an LPLR, this ablation suggests that width can also influence which favorable region is reached (i.e., the loss level of the attained solution), thereby linking over-parameterization to improved optimization outcomes in our setting (see Table 1 and Figure 5).
+7 Discussion and Conclusion
+This paper addresses a central question in deep learning theory: why gradient-based methods often converge rapidly in practice despite the highly non-convex nature of neural network loss landscapes.
+Building on the Locally Quasi-Convex Regions (LQCRs) framework (Aich et al., 2025), we show that the same local geometry that guarantees monotonic progress can, under a mild NTK stability assumption restricted to the local region of interest (Jacot et al., 2018), also imply a local Polyak-Łojasiewicz (PL)-type inequality (Polyak, 1963).
+This provides a principled route from previously established sublinear guarantees to locally linear convergence behavior consistent with empirical training dynamics.
+Our perspective complements recent advances.
+Architectural approaches such as PLNet enforce PL-like conditions by design (Wang et al., 2024), while we aim to explain why favorable optimization behavior can arise in standard architectures without modification.
+Stochastic analyses of SGD under local Łojasiewicz-type conditions (An and Lu, 2024) yield important rate statements but typically assume stronger local regularity; our geometric viewpoint provides a concrete mechanism for why such conditions may hold locally along training trajectories.
+Related progress for linear networks establishes local PL behavior under relaxed assumptions (Xu et al., 2025); our work extends the conceptual thread to nonlinear, finite-width networks.
+Our main theoretical contribution is the introduction and analysis of the Locally Polyak–Łojasiewicz Region (LPLR), formalizing that well-behaved local regions can support fast rates through a PL-like error bound.
+The analysis is stated for finite-width networks under local conditions, with the NTK serving as the bridge between geometry and optimization via the stability assumption (Jacot et al., 2018).
+Conceptually, this links local landscape structure (via LQCRs) (Aich et al., 2025), kernel-based linearization intuition (via the NTK) (Jacot et al., 2018), and optimization theory converting PL-type inequalities into linear-rate guarantees (Polyak, 1963).
+Empirically, the results are consistent with the practical relevance of the LPLR framework across controlled and more realistic settings.
+In a full-batch MNIST experiment, the observed loss dynamics exhibit linear-rate decay and a PL-like relationship between the squared gradient norm and suboptimality over the explored region.
+Initialization also substantially affects convergence speed, underscoring the importance of the local region entered early in training.
+In a ResNet-style CNN trained on a CIFAR-10 subset with mini-batch SGD and cosine annealing, the same qualitative signatures remain visible after smoothing, suggesting that similar behavior can persist under modern architectures and stochastic optimization.
+From a practitioner’s standpoint, the framework suggests that (i) initialization can influence the effective rate constant by determining which local region training enters, (ii) design choices that improve conditioning and stabilize training may indirectly strengthen local PL-like behavior by promoting local NTK stability (Jacot et al., 2018), and (iii) relatively large learning rates can still yield fast progress when compatible with smoothness constraints and when iterates remain within the favorable region.
+We also note limitations and directions for future work.
+Our primary theory is for full-batch GD; a rigorous stochastic extension would need to quantify convergence to a noise ball and characterize how stochasticity interacts with remaining within an LPLR, building on analyses under PL/KL-type conditions (Karimi et al., 2016; Kingma and Ba, 2015; Reddi et al., 2018).
+Moreover, while local NTK stability is supported empirically in our setting, understanding when it holds and how it scales with width, depth, normalization, and data complexity remains important.
+Finally, connecting LPLR properties to generalization and implicit regularization is a promising direction.
+In conclusion, Locally Polyak–Łojasiewicz Regions provide a mechanism linking local landscape geometry to fast convergence rates in neural network training.
+By connecting LQCR-type geometric stability (Aich et al., 2025) with local NTK stability (Jacot et al., 2018) and PL-based optimization theory (Polyak, 1963), the framework offers theoretical clarity and empirically supported intuition for rapid convergence in realistic finite-width regimes.
+Impact Statement
+This work advances the theoretical foundations of deep learning by providing a principled explanation for a widely observed empirical phenomenon: the rapid convergence of gradient-based optimization in highly non-convex neural network training.
+By connecting local geometric structure and NTK stability to a local PL-type mechanism, our results can help researchers reason more precisely about when fast training should be expected and how design choices (e.g., width, normalization, initialization, and learning-rate schedules) may influence optimization behavior.
+While the paper is theoretical, clearer guarantees and diagnostic tools for optimization can support more reliable and resource-efficient model development, reducing trial-and-error and potentially lowering the computational cost and environmental footprint of training.
+We do not anticipate direct negative societal impacts from this work beyond the standard dual-use considerations that apply broadly to deep learning methods.